Problem 13-5A Suppose selected financial data of Target and Wal-Mart for 2017 are presented here (in millions). Wal-Mart Target Income Statement Data for Net sales Cost of goods sold Selling and administrative expenses Interest expense Other income (expense) Income tax expense Net income $66,900 45,000 15,000 $419,000 309,000 80,000 1,900 (80) 1,500 (390) 6,700 4,620 21,010 Balance Sheet Data (End of Year) Current assets Noncurrent assets Total assets Current liabilities Long-term debt Total stockholders equity Total liabilities and stockholders $17,000 26,700 $43,700 $11,000 18,200 14,500 $43,700 $45,000 120,000 $165,000 55,000 45,000 65,000 $165,000 Total assets Total stockholders equity Current liabilities Total labilities Beginning-of-Year Balances $165,000 64,000 58,000 101,000 $43,000 12,600 30,400 2 4 6 8 9
Other Data Average net accounts receivable Average inventory Net cash provided by operating $7,700 6,800 S,500 1,800 490 $4,200 34,500 25,800 12,300 4,200 activities Capital expenditures Dividends (a) For each company, compute the following ratios. (Round all answers to 2 decimal places, eg. 1.83 or 1.83%.) Ratio Target Wal-Mart (1) Current ratio (2) Accounts receivable turnover (3) Average collection period (4) Inventory turnover (5) Days in inventory (6) Profit margin (7) Asset turnover (8) Return on assets (9) Return on common stockholders equity (10) Debt to assets ratio (11) Times interest earned (12) Free cash fow Click if you would like to Show Work for this question: times days days times days times Open Show Work 8 2
Answer #1
Target Corporation Walmart Stores, Onc.
1 Current ratio = Current Assets/ Current Liabilities 1.7 0.78 Times
2 Receivables turnover (Times per year) = Net sales/ Av. Net Receivables 8.69 99.76 Times
3 Average collection period (Days) = 365/Receivable turnover 42.01 3.66 Days
4 Inventory turnover (Times per year) = Cost of Good Sold/ Av.Inventory 6.62 8.96 Times
5 Days in Inventory (Days) 55.16 40.75 Days
6 Profit margin (%) = Net Income/ Net sales 6.91% 5.01%
7 Asset turnover = Net Sales/ Av. Total Assets 1.54 2.54 Times
8 Return on assets (%) = Net Profit/ Av. Total Assets 10.66% 12.73%
9 Return on common stockholders' equity (%) = Net Profit/ Av. ShareholderEquity 34.10% 32.57%
10 Debt to total assets (%) = Average total Debt/ Average Total Assets 68.74% 60.91%
11 Times interest earned = net income + interest expense + income tax / Interest Expenses 9.74 15.58 Times
Target Corporation Walmart Stores, Onc.
1 Current ratio = Current Assets/ Current Liabilities =B11/B21 =C11/C21 Times
2 Receivables turnover (Times per year) = Net sales/ Av. Net Receivables =B3/B24 =C3/C24 Times
3 Average collection period (Days) = 365/Receivable turnover =365/C31 =365/D31 Days
4 Inventory turnover (Times per year) = Cost of Good Sold/ Av.Inventory =B4/B25 =C4/C25 Times
5 Days in Inventory (Days) =365/C33 =365/D33 Days
6 Profit margin (%) = Net Income/ Net sales =B9/B3 =C9/C3
7 Asset turnover = Net Sales/ Av. Total Assets =B3/AVERAGE(B13,B19) =C3/AVERAGE(C13,C19) Times
8 Return on assets (%) = Net Profit/ Av. Total Assets =B9/AVERAGE(B13,B19) =C9/AVERAGE(C13,C19)
9 Return on common stockholders' equity (%) = Net Profit/ Av. ShareholderEquity =B9/AVERAGE(B16,B20) =C9/AVERAGE(C16,C20)
10 Debt to total assets (%) = Average total Debt/ Average Total Assets =AVERAGE((B14+B15),B22)/AVERAGE(B13,B19) =AVERAGE((C14+C15),C22)/AVERAGE(C13,C19)
11 Times interest earned = net income + interest expense + income tax / Interest Expenses =(B9+B6+B8)/B6 =(C9+C6+C8)/C6 Times


Income Statement data for year Wal-Mart Stores,Inc. Target 3 Net sales 4 Cost of goods sold 5 Selling and administrati 6 Interest expense 7 Other income (expense) 8 Income tax expense 9 Net Income 10 11 Current assets 12 noncurrent assets 13 total assets 14 current liabilities 15 long-term debt 16 total stockholders equit 17 total liabilites and stockl 18 19 total assets 20 total stockholders equit 21 current liabilities 22 total liabilities 23 24 Average net receivables 25 average inventory 26 net cash provided by ope 27 Capital expenditure 28 Dividend 66,9004,19,000 5,0003,09,000 15,00080,000 1,900 390 6,700 21,010 700 80 4,620 Balance sheet data (end of year) 17,000 45,000 26,7001,20,000 43,700 1,65,000 1,000 55,000 18,200 45,000 4,500 65,000 43,700 1,65,000 Beginning-of-year balances 430001,65,000 64,000 58,000 30,400 1,01,000 12,600 10,000 Other Data 7,700 4,200 6,80034,500 5,50025,800 12,300 4,200 490